OpenStudy (volleyballlover55):

Compare 8 ⋅ 10^7 to 2 ⋅ 10^4. 8 ⋅ 10^7 is 4,000 times larger than 2 ⋅ 10^4. 8 ⋅ 10^7 is 400 times larger than 2 ⋅ 10^4. 8 ⋅ 10^7 is 40 times larger than 2 ⋅ 10^4. 8 ⋅ 10^7 is 4 times larger than 2 ⋅ 10^4.

OpenStudy (volleyballlover55):

@greatlife44

OpenStudy (volleyballlover55):

@rebeccaxhawaii

OpenStudy (anonymous):

The answer is 4,000

OpenStudy (anonymous):

8 * 10^7 = 80,000,000 2 * 10^4 = 20,000 20,000 * 4000 = 800,000

rebeccaxhawaii (rebeccaxhawaii):

got it?